OpenStudy (anonymous):

How do you simplify B= log (I/ 10^-16) ?

OpenStudy (raden):

hint : 1/10^-6 = 10^6

OpenStudy (anonymous):

Umm, okay im gonna change I to x so B= log ( x/ 10^-16) would it be log x - log 10 ^ -16 ?

OpenStudy (raden):

no, that should be log x + log 10^6 remmber that : log (a * b) = log a + log b

OpenStudy (anonymous):

but...... its log (x DIVIDED BY 10^-16)

OpenStudy (anonymous):

\[\log \frac{ x }{ 10^{-16} }\]

OpenStudy (anonymous):

so it'd be subtraction npt addition